OpenStudy (anonymous):

Based on the lesson, identify the first operation needed to solve the following equation. x/3 - 3 = 11 A addition B subtraction C multiplication D division

OpenStudy (anonymous):

@tywower

OpenStudy (anonymous):

@ImaGuest

OpenStudy (jdoe0001):

use \(\huge P\quad E\quad M\quad D\quad A\quad S\)

OpenStudy (jdoe0001):

hmmm actaully, this is just a linear equation hmmm

OpenStudy (jdoe0001):

to isolate "x" all by herself on the left-side firstly, get rid of the "-3" then get rid of the 3 in the denominator

OpenStudy (anonymous):

i dont feel like it @jdoe0001

OpenStudy (jdoe0001):

anyhow \(\bf \cfrac{x}{3}\cancel{-3{\color{brown}{ +3}}}=11{\color{brown}{ +3}}\implies \cfrac{x}{3}=11+3 \\ \quad \\ {\color{brown}{ \cancel{3}}}\cdot \cfrac{x}{\cancel{3}}={\color{brown}{ 3}}\cdot (11+3)\)